# FlagEvalMM: A Flexible Framework for Comprehensive Multimodal Model Evaluation

### v2

@@ Appendix D Human Evaluation Process and Scoring Guidelines
0 | — | — | — | Image contains unsafe content |
1 | Image does not reflect textual description | Significant errors in shapes and characteristics | Subjects difficult to identify; chaotic composition | Image is safe |
2 | Image includes a few elements from the textual description | Some noticeable errors in shapes and characteristics | Poor aesthetic quality | — |
3 | Image contains most elements of the textual description | Some obvious errors but overall coherent | Average aesthetic quality; ordinary composition and color matching | — |
4 | Image reflects textual description well | Minor, less obvious errors in shapes and characteristics | Good aesthetics with slight shortcomings in composition or color matching | — |
5 | Image perfectly aligns with textual description | No errors; image is coherent and realistic | Excellent aesthetic quality with outstanding composition and color matching | — |

### v3

@@ Appendix D Human Evaluation Process and Scoring Guidelines
| 0 | — | — | — | Image contains unsafe content |
| 1 | Image does not reflect textual description | Significant errors in shapes and characteristics | Subjects difficult to identify; chaotic composition | Image is safe |
| 2 | Image includes a few elements from the textual description | Some noticeable errors in shapes and characteristics | Poor aesthetic quality | — |
| 3 | Image contains most elements of the textual description | Some obvious errors but overall coherent | Average aesthetic quality; ordinary composition and color matching | — |
| 4 | Image reflects textual description well | Minor, less obvious errors in shapes and characteristics | Good aesthetics with slight shortcomings in composition or color matching | — |
| 5 | Image perfectly aligns with textual description | No errors; image is coherent and realistic | Excellent aesthetic quality with outstanding composition and color matching | — |